Calculate the value of the ionization constant Ka, if the pKa is 5.64


1
Expert's answer
2021-08-23T02:56:20-0400

Solution:

pKa = - log(Ka)

Ka = 10-pKa

Ka = 10-5.64 = 2.29×10-6

Ka = 2.29×10-6


Answer: The value of the ionization constant (Ka) is 2.29×10-6